OpenStudy (anonymous):

10xy^2/3z divided by 5xy/6z^3

OpenStudy (anonymous):

I'll reduce it for you: 10/3 divided by 5/6 :)

OpenStudy (anonymous):

@kels200105 how is that helping? How is that telling the person how to find the answer? I know how to answer this question and I have no clue what you did there.

OpenStudy (anonymous):

4yz^2 is the answer

OpenStudy (anonymous):

\[\frac{ 10xy^2 }{ 3z } \times \frac{ 6z^3 }{ 5xy }\]

OpenStudy (anonymous):

Cross simplify

OpenStudy (anonymous):

\[\frac{ (2y)(2z^2) }{ 1 }\]